Money demand in an economy in which no interest is paid on money is Md/P =...

  1. Money demand in an economy in which no interest is paid on money is

Md/P = 5000 + 0.2Y + 1000i.

You know that P = 100, Y = 1000, and i = 0.10. Find

  1. real money demand.

  1. nominal money demand.

  1. money velocity.
Answer #1

#ma = 5000 +0.2 Y+1000 i P2 1000 Y2 1000 ia 0.01 @ Real Money Demand me 5000 +0.27 + 1000 i = 5000 + (0.2x1000) + (10008 0.01③ According Quantity theory of Money MO2 px IM:- Nominal Money Demand vi vilocity Pir Price Y:- output 530000 X 1 z 100 x 100